# Can Language Models Verbalize Well-Defined Notions of Uncertainty?

### v2

@@ Abstract
As large language models (LLMs) are increasingly used for factual question-answering, it becomes more important for LLMs to have the capability to communicate the likelihood that their answer is correct. For these verbalized expressions of uncertainty to be meaningful, they should reflect the error rates at the expressed level of confidence. However, when prompted to express confidence, the error rates of current LLMs are inconsistent with their communicated confidences, highlighting the need for uncertainty quantification methods. Many prior methods calculate lexical uncertainty, estimating a model’s confidence in the specific string it generated. In some cases, however, it may be more useful to estimate semantic uncertainty, or the model’s confidence in the answer regardless of how it is verbalized. We propose a simple procedure, uncertainty distillation, to teach an LLM to verbalize calibrated semantic confidences. Using held-out data to map initial uncertainty estimates to meaningful probabilities, we create examples annotated with verbalized probabilities for supervised fine-tuning. We compare uncertainty distillation to several strong baselines, and find that our method yields verbalized confidences that correlate well with observed error rates.
Uncertainty Distillation: Teaching Language Models to Express Semantic Confidence
Sophia Hager, David Mueller, Kevin Duh, Nicholas Andrews Johns Hopkins University {shager2,noa}@jhu.edu
@@ 1 Introduction
Advances in LLM research have led to instruction-tuned generative models with impressive capabilities on many challenging tasks (OpenAI et al., 2024; Jiang et al., 2023; Dubey et al., 2024). While the flexibility and quality of these models is appealing, they may still hallucinate or give incorrect answers (Rawte et al., 2023; Bai et al., 2024). However, language models do not readily provide an interpretable measure of a model’s likelihood of correctness. LLMs tend to produce poorly-calibrated confidences when prompted to do so, and are often confidently incorrect Xiong et al. (2024). Furthermore, the elicited confidences may be impacted in unexpected ways by the choice of prompt (Sclar et al., 2023), such as the interpretation of “very confident” being dependent on the wording of the prompt.
@@
There are several other approaches as an alternative to prompting. Models’ token-level probabilities can be used to provide information as a measure of lexical uncertainty, which gives information about the likelihood of a generated string. This is often useful; however, the same fact can be expressed in any number of ways—“Berlin’s the capital of Germany” or “The capital of Germany is Berlin!” or “Die Hauptstadt Deutschlands ist Berlin”—all capturing the same meaning Kuhn et al. (2023). Semantic uncertainty is therefore challenging to capture, as token-level probabilities are influenced by the phrasing of an answer just as much as the semantics of the answer itself. This issue is particularly challenging for models employing large vocabularies such as multilingual language models, language models employing byte or character-level tokenization, or when using LLMs that are prone to producing extraneous outputs Xue et al. (2021); Wang et al. (2024).
@@
Figure 1: An overview of how our method might be applied to a simple QA example. In Monte Carlo sampling with semantic normalization, we sample repeatedly from our language model, and use a normalization function to consolidate answers with the same semantic meaning. By consolidating the counts, we obtain a Monte Carlo estimate of each answer’s probability. In post-hoc calibration, we pass this estimate through a learned post-hoc calibration function to better align it with its likelihood of correctness. Finally, in self-annotation and fine-tuning, we translate these probabilities to verbalized signifiers and fine-tune a model to output verbalized confidences in addition to the answer.
@@
We present uncertainty distillation, a scheme for fine-tuning a language model to verbalize uncertainty based on its own internal state. Notably, uncertainty distillation teaches models to estimate their semantic—rather than lexical—uncertainty, as the distilled confidences are estimated from the probabilities of semantically normalized outputs, rather than relying on token-level probabilities. At inference time, models trained using uncertainty distillation efficiently generate a well-calibrated and interpretable statement of confidence in their answers, such as “Berlin is the capital of Germany [high confidence].” [^1] The uncertainty could be expressed in a variety of ways, including using special characters or numeric values. Our approach enables semantically equivalent but lexically different predictions to be assigned the same confidence, and a single generation with multiple claims can each be assigned different confidences. Additionally, uncertainty distillation is computationally inexpensive at inference time, generating only a handful of additional tokens. And compared to methods such as P(IK) Farquhar et al. (2024), we do not require a separate uncertainty network; our approach uses standard supervised fine-tuning recipes for LLMs.
@@
Rather than specializing to a specific model or task, uncertainty distillation involves self-annotation of any desired QA dataset with the base model’s calibrated uncertainties, which are then used to fine-tune that model to produce verbalized confidences. At a high level (Figure 1), our approach consists of three steps: (1) obtaining semantic uncertainty estimates from the model on held-out calibration data; (2) post-hoc calibrating these into meaningful probabilities; and (3) teaching the model via supervised fine-tuning to output verbalized confidences along with its predictions.
@@
  • We propose uncertainty distillation, a simple yet effective scheme which uses supervised fine-tuning to teach LLMs to output calibrated semantic confidence statements along with their predictions.
@@
  • We analyze the impact of choices made when implementing uncertainty distillation.
@@ 2 Related Work
Generally, calibration refers to the concept that predicted probabilities should align with the probability of correctness (Guo et al., 2017). Mielke et al. (2022) additionally propose the conception of “linguistic calibration"—that models demonstrate uncertainty or doubt through natural language when they are incorrect, determining this uncertainty by using a predictor to determine the likelihood that an answer is correct and considering that to be the model’s uncertainty. There are significant advantages to verbalizing uncertainty: for one, there is relatively low computational overhead to generate several extra tokens, while using a separate calibration model to estimate confidence and then communicate this information to the user requires more computation at inference time Yang et al. (2024). Verbalized confidences are also readily interpretable to an LLM when reasoning about uncertainty, or to an average end-user regardless of experience or background.
@@
Lexical uncertainty quantification metrics using information from token-level probabilities are commonly used and frequently effective Hu et al. (2023); Malinin and Gales (2021). These probabilities are easily obtainable, do not require additional inference-time compute to generate, and often provide sufficient information for downstream use cases: e.g. error correction in chain of thought Yin et al. (2024), hallucination detection Arteaga et al. (2024), or out-of-distribution data detection Hendrycks et al. (2020b). However, there are several disadvantages to lexical uncertainty quantification: it relies on model probabilities which may not be well-calibrated Guo et al. (2017), and is often ineffective on calculating uncertainty of long generations Zhang et al. (2024). The latter, in particular, may present problems for end users, as models trained using Reinforcement Learning from Human Feedback (RLHF) are often incentivized to produce long outputs Singhal et al. (2024). It is therefore important to consider uncertainty quantification methods that do not rely on token-level probabilities to estimate uncertainty.
@@
In contexts where lexical uncertainty falls short, a natural method to obtain verbalized confidences might be to simply prompt a model to output confidences. This would theoretically provide an estimate of uncertainty without explicitly using token-level probabilities to estimate confidence. However, in practice, LLMs tend to overestimate their own confidence, possibly because human annotators tend to prefer texts with fewer markers of uncertainty Zhou et al. (2024). This, in turn, suggests while simply altering prompts may result in improved confidence estimates (Xiong et al., 2024; Tian et al., 2023), models may be fundamentally limited in their ability to acknowledge uncertainty without further training.
Running multiple steps at inference time may provide a better estimate of semantic probability. Xiong et al. (2024) investigate several inference-time strategies which use multiple steps to estimate model uncertainty, such as sampling several answers on the same question or noting if a model changes its answer when prompted with a misleading alternative. While these methods do lead to improvements in LLM calibration, no single intervention consistently emerges as the most successful, and the authors note there is significant scope for improvement. Kuhn et al. (2023) and Farquhar et al. (2024) more explicitly relate this to semantic uncertainty, and find that sampling $m$ predictions from the model and clustering by semantic equivalence results in a robust measure of semantic uncertainty that compares favorably to lexical uncertainty. A major disadvantage of these sampling-based approaches is their increased computational complexity at inference time, however; for instance, the semantic clustering approach of Farquhar et al. (2024), which we compare to in our experiments, requires 20 samples at inference time and calls to a separate entailment model.
@@ 3.1 Monte Carlo sampling with semantic normalization
Assuming input $x$ and output $y$, we are looking to find $P(y\in Y_{\text{equivalent}}\mid x)$, the model’s likelihood of producing this answer or one that is semantically equivalent; however, this would require marginalization over an infinite set of strings $Y$. To make this a tractable problem, we use a Monte Carlo approximation, where our estimate of the models’ predictive distribution improves with $N$, at the expense of additional computational cost. Note however that we do not assume this quantity is a meaningful probability out-of-the-box due to potential overfitting or underfitting of the base model. To diagnose potential miscalibration of the base model as well as correct for it, we may use held-out calibration data that was not seen during training. [^2] We examine the efficacy of our method when this assumption does not hold true in §missing 6.1.
In more detail, to fit a post-hoc calibrator, we need a supervised dataset of datapoints not seen at training time $\{X^{\text{cal}},Y^{\text{cal}}\}$. For each example $x\in X^{\text{cal}}$ we sample $N$ candidate answers $\{\hat{y}_{i}\}_{i=1}^{N}\sim P_{\theta}(Y\mid X=x)$ from a model’s predictive distribution [^3] This model may have been fine-tuned on the specific task as in Appendix A or instruction-tuned as in §missing 4.2.. Before calculating the relative frequency of strings, we apply a normalization function (or set of normalization functions) to consolidate semantically similar outputs. In the short-form QA tasks we consider in §missing 4, we use the simple normalization function of isolating a multiple choice answer using tags, removing punctuation and standardizing capitalization, but for more complex questions other approaches may be required (Huang et al., 2024). For example, Kuhn et al. (2023) use natural language inference to cluster semantically equivalent answers and Tian et al. (2023) use an LLM as a judge of correctness. After consolidating similar events, the relative frequency $f$ of these events is a measure of the LLM’s uncertainty in those events, although this may not be a well-calibrated probability.
@@ 3.2 Post-hoc calibration
In general, neural networks are prone to miscalibration, even when trained using proper losses such as cross-entropy. A common remedy is to apply post-hoc calibration methods, which usually involve some form of regression on the predicted scores to transform them into meaningful probabilities. Specifically, we post-hoc calibrate the relative frequencies of each semantic cluster found in the previous step. Two common options for post-hoc calibration are isotonic regression and Platt scaling (sometimes called temperature scaling) (Guo et al., 2017). Our approach uses a model’s predictions on $\{X^{\text{cal}},Y^{\text{cal}}\}$ to diagnose and mitigate badly-calibrated initial model probabilities. We fit an isotonic regression model [^4] Implemented using scikit-learn 1.5.2 on our calibration set by comparing the predicted scores to observed labels. [^5] We discuss post-hoc calibration further in Appendix D. We compare each prediction $\hat{y}$ with score $f$ to observed events $y$. This yields a calibration map $c:\mathbb{R}\rightarrow[0,1]$ we apply to the relative frequencies of events from samples in the previous step to yield probabilities.
@@ 3.3 Self-annotation and fine-tuning
We compute the calibrated probability $p=c(f)$ associated with each prediction in the held-out calibration data, and choose a mapping into discrete confidence bins. Several options are possible for this binning function $b$, including adaptive schemes as well as uniform schemes, the number of bins $B$, and so on. In this paper, we focus on a simple fixed-width scheme with $5$ bins. Let $\hat{Y}$ denote the set of all predictions on $X^{\text{cal}}$, and, if the model was previously fine-tuned on a supervised training set $X^{\text{train}}$, we include predictions on $X^{\text{train}}$. We deterministically transform each prediction and calibrated confidence into a training example for a round of supervised fine-tuning by verbalizing the corresponding bin in the answer. For example, the fifth of five bins may correspond to “very high confidence.” The token sequences chosen to encode each bin are arbitrary; for easy interpretability, we use short descriptors in this paper, namely “very low confidence,” “low confidence,” “medium confidence,” “high confidence,” and “very high confidence.”
@@
While correct answers should always be added as training data, appending the confidence scores to incorrect answers may improve the model’s ability to correctly verbalize its own confidence. However, it may also decrease the accuracy of the QA model. We introduce a hyperparameter to control the number of incorrect answers added to the training data. In §missing 6.2, we further investigate the impact of this hyperparameter.
Starting from the sampled model, we perform supervised fine-tuning on these self-annotated targets with verbalized confidences to estimate a second model capable of verbalizing its confidence. If training an instruction-tuned model, we append an additional instruction such as "Additionally state how confident you are in your answer." to the preexisting instruction [^6] See Appendix C for details on the specific prompts used in each experiment.. If a reasoning trace has been generated during sampling, we randomly select a reasoning trace to add to the target answer from all possible options. At inference time, we obtain predictions and verbalized confidences from this new model on held-out test data. We remark that our model incurs little additional cost at inference time, as opposed to other confidence elicitation methods which require inference-time sampling Farquhar et al. (2024); Xiong et al. (2024).

### v3

# Uncertainty Distillation: Teaching Language Models to Express Semantic Confidence

@@ Abstract
As large language models (LLMs) are increasingly used for factual question-answering, it becomes more important for LLMs to have the capability to communicate the likelihood that their answer is correct. For these verbalized expressions of uncertainty to be meaningful, they should reflect the error rates at the expressed level of confidence. However, when prompted to express confidence, the error rates of current LLMs are inconsistent with their communicated confidences, highlighting the need for uncertainty quantification methods. Many prior methods calculate lexical uncertainty, estimating a model’s confidence in the specific string it generated. In some cases, however, it may be more useful to estimate semantic uncertainty, or the model’s confidence in the answer regardless of how it is verbalized. We propose a simple procedure, uncertainty distillation, to teach an LLM to verbalize calibrated semantic confidences. Using held-out data to map initial uncertainty estimates to meaningful probabilities, we create examples annotated with verbalized probabilities for supervised fine-tuning. We find that our method yields verbalized confidences that correlate well with observed error rates, even when compared to strong baselines, some of which are more than twenty times slower at inference time. Additionally, we demonstrate that our method can be applied to black-box models that allow API-based fine-tuning, resulting in estimates of uncertainty that are both more effective and more efficient than any of our baselines.
@@ 1 Introduction
Advances in LLM research have led to instruction-tuned generative models with impressive capabilities on many challenging tasks (openai2024gpt4technicalreport; jiang2023mistral7b; dubey2024llama3herdmodels). While the flexibility and quality of these models is appealing, they may still hallucinate or give incorrect answers (rawte-etal-2023-troubling; Bai2024HallucinationOM). However, language models do not readily provide an interpretable measure of a model’s likelihood of correctness. LLMs tend to produce poorly-calibrated confidences when prompted to do so, and are often confidently incorrect (xiong2024can). Furthermore, the elicited confidences may be impacted in unexpected ways by the choice of prompt (sclar2023quantifying), such as the interpretation of “very confident” being dependent on the wording of the prompt.
@@
There are several other approaches as an alternative to prompting. Models’ token-level probabilities can be used to provide information as a measure of lexical uncertainty, which gives information about the likelihood of a generated string. This is often useful; however, the same fact can be expressed in any number of ways—“Berlin’s the capital of Germany” or “The capital of Germany is Berlin!” or “Die Hauptstadt Deutschlands ist Berlin”—all capturing the same meaning (kuhn2023semantic). Semantic uncertainty is therefore challenging to capture, as token-level probabilities are influenced by the phrasing of an answer just as much as the semantics of the answer itself. This issue is particularly challenging for models employing large vocabularies such as multilingual language models, language models employing byte or character-level tokenization, or when using LLMs that are prone to producing extraneous outputs (xue2021mt5massivelymultilingualpretrained; wang2024mambabyte).
@@
Figure 1: An overview of our method, Uncertainty Distillation. At training time, in Monte Carlo sampling with semantic normalization, we sample repeatedly from our language model, and use a normalization function to consolidate answers with the same semantic meaning. By consolidating the counts, we obtain a Monte Carlo estimate of each answer’s probability. In post-hoc calibration, we pass this estimate through a learned post-hoc calibration function to better align it with its likelihood of correctness. Finally, in self-annotation and fine-tuning, we translate these probabilities to verbalized signifiers and fine-tune a model to output verbalized confidences in addition to the answer. This method confers several advantages, listed in the table: at inference time, a single model generates the confidence efficiently in a single pass, providing high discriminative power with little computational overhead. The length of the answer does not directly impact the confidence, and white-box access to weights is not required.
@@
We present uncertainty distillation [^1] We choose this name to evoke model distillation, a process which like uncertainty distillation requires an offline cost to generate data to train a more efficient model., a scheme for fine-tuning a language model to verbalize uncertainty based on its own internal state. Notably, uncertainty distillation teaches models to estimate their semantic—rather than lexical—uncertainty, as the distilled confidences are estimated from the probabilities of semantically normalized outputs, rather than relying on token-level probabilities. At inference time, models trained using uncertainty distillation efficiently generate a well-calibrated and interpretable statement of confidence in their answers, such as “Berlin is the capital of Germany [high confidence].” [^2] The uncertainty could be expressed in a variety of ways, including using special characters or numeric values. Our approach enables semantically equivalent but lexically different predictions to be assigned the same confidence, and a single generation with multiple claims can each be assigned different confidences. Uncertainty distillation is computationally inexpensive at inference time, generating only a handful of additional tokens. Compared to methods such as P(IK) (Farquhar2024), we do not require a separate uncertainty network; our approach uses standard supervised fine-tuning recipes for LLMs. Our method can be applied to open-source LLMs as well as proprietary LLMs that allow fine-tuning; white-box access to model weights is not required.
@@
Uncertainty distillation involves self-annotation of any desired QA dataset with the base model’s calibrated uncertainties, which are then used to fine-tune that model to produce verbalized confidences. At a high level (Figure 1), our approach consists of three steps: (1) obtaining semantic uncertainty estimates from the model; (2) post-hoc calibrating these into meaningful probabilities; and (3) teaching the model via supervised fine-tuning to output verbalized confidences along with its predictions.
@@
  • We propose uncertainty distillation, a simple yet effective scheme which uses supervised fine-tuning to teach LLMs to output calibrated semantic confidence statements along with their predictions. We publish our code and trained models. [^3] https://anonymous.4open.science/r/uncertainty-distillation-anon-05CB/README.md contains necessary code to replicate the results for API tuning.
@@
  • We analyze whether models trained with uncertainty distillation can apply their representations of uncertainty to unseen topics at inference time without further fine-tuning.
@@ 2 Related Work
Generally, calibration refers to the concept that predicted probabilities should align with the probability of correctness (Guo-etal-calibration-2017). mielke-etal-2022-reducing additionally propose the conception of “linguistic calibration”—that models demonstrate uncertainty or doubt through natural language when they are incorrect, determining this uncertainty by using a predictor to determine the likelihood that an answer is correct and considering that to be the model’s uncertainty. There are significant advantages to verbalizing uncertainty: for one, there is relatively low computational overhead to generate several extra tokens, while using a separate calibration model to estimate confidence and then communicate this information to the user requires more computation at inference time (yang2024verbalizedconfidencescoresllms). Verbalized confidences are also readily interpretable to an LLM when reasoning about uncertainty, or to an average end-user regardless of experience or background.
@@
Lexical uncertainty quantification metrics using information from token-level probabilities are commonly used and frequently effective (hu2023uncertaintynaturallanguageprocessing; malinin2021uncertaintyestimationautoregressivestructured). These probabilities are easily obtainable, do not require additional inference-time compute to generate, and often provide sufficient information for downstream use cases: e.g. error correction in chain of thought (yin-etal-2024-reasoning), hallucination detection (arteaga2024hallucination), or out-of-distribution data detection (hendrycks2020pretrainedtransformersimproveoutofdistribution). However, there are several disadvantages to lexical uncertainty quantification: it relies on model probabilities which may not be well-calibrated (Guo-etal-calibration-2017), and is often ineffective on calculating uncertainty of long generations (zhang2024luqlongtextuncertaintyquantification). The latter, in particular, may present problems for end users, as models trained using Reinforcement Learning from Human Feedback (RLHF) are often incentivized to produce long outputs (singhal2024a). It is therefore important to consider uncertainty quantification methods that do not rely on token-level probabilities to estimate uncertainty.
@@
In contexts where lexical uncertainty falls short, a natural method to obtain verbalized confidences might be to simply prompt a model to output confidences, providing an estimate of uncertainty without explicitly using token-level probabilities. However, in practice, LLMs tend to overestimate their own confidence, possibly because human annotators tend to prefer texts with fewer markers of uncertainty (zhou-etal-2024-relying). This, in turn, suggests while simply altering prompts may result in improved confidence estimates (xiong2024can; tian-etal-2023-just), models may be fundamentally limited in their ability to acknowledge uncertainty without further training.
Running multiple steps at inference time may provide a better estimate of semantic probability. xiong2024can investigate several inference-time strategies which use multiple steps to estimate model uncertainty, such as sampling several answers on the same question or noting if a model changes its answer when prompted with a misleading alternative. While these methods do lead to improvements in LLM calibration, no single intervention consistently emerges as the most successful, and the authors note there is significant scope for improvement. kuhn2023semantic and Farquhar2024 more explicitly relate this to semantic uncertainty, and find that sampling $m$ predictions from the model and clustering by semantic equivalence results in a robust measure of semantic uncertainty that compares favorably to lexical uncertainty. A major disadvantage of these sampling-based approaches is their increased computational complexity at inference time, however; for instance, the semantic clustering approach of Farquhar2024, which we compare to in our experiments, requires 20 samples and calls to a separate entailment model at inference time.
@@ 3.1 Monte Carlo sampling with semantic normalization
Assuming input $x$ and output $y$, we are looking to find $\sum_{y\in Y_{\text{equivalent}}}P(y\mid x)$, the model’s likelihood of producing this answer or one that is semantically equivalent; however this would require marginalization over an infinite set of strings $Y$. To make this a tractable problem, we use a Monte Carlo approximation, where our estimate of the models’ predictive distribution improves with $N$, at the expense of additional offline computation. Note however that we do not assume this quantity is a meaningful probability out-of-the-box due to potential overfitting or underfitting of the base model. To diagnose potential miscalibration of the base model as well as correct for it, we may fit a post-hoc calibrator if the training data demonstrates miscalibration.
In more detail, to fit a post-hoc calibrator, we need a supervised dataset of datapoints not seen at training time $\{X^{\text{cal}},Y^{\text{cal}}\}$. For each example $x\in X^{\text{cal}}$ we sample $N$ candidate answers $\{\hat{y}_{i}\}_{i=1}^{N}\sim P_{\theta}(Y\mid X=x)$ from a model’s predictive distribution [^4] This model may have been fine-tuned on the specific task as in Appendix C or instruction-tuned as in § 4 andAppendix B.. Before calculating the relative frequency of strings, we apply a normalization function (or set of normalization functions) to consolidate semantically similar outputs. In the short-form QA tasks we consider in § 4, we use the simple normalization function of isolating a multiple choice answer using tags, removing punctuation and standardizing capitalization; we demonstrate how semantic normalization can be applied to more complex tasks in Appendix A. After consolidating strings belonging to the same event, the relative frequency $f$ of these events is a measure of the LLM’s uncertainty in those events, although this may not be a well-calibrated probability.
@@ 3.2 Post-hoc calibration
Neural networks are prone to miscalibration. A common remedy is to apply post-hoc calibration methods, which usually involve some form of regression on predicted scores to transform them into meaningful probabilities. Specifically, we post-hoc calibrate the relative frequencies of each semantic cluster found in the previous step. Two common options for post-hoc calibration are isotonic regression and Platt scaling (sometimes called temperature scaling) (Guo-etal-calibration-2017). Our approach uses a model’s predictions on $\{X^{\text{cal}},Y^{\text{cal}}\}$ to diagnose and mitigate badly-calibrated initial model probabilities. We fit an isotonic regression model [^5] We use isotonic regression for ease of training and use; this could be replaced with a different post-hoc calibration method, or omitted entirely as discussed in Appendix G. We use the scikit-learn 1.5.2 with no modification. on our calibration set by comparing the predicted scores to observed labels. [^6] We discuss the effect of post-hoc calibration further in Appendix G. We compare each prediction $\hat{y}$ with score $f$ to observed events $y$. This yields a calibration map $c:\mathbb{R}\rightarrow[0,1]$ we apply to the relative frequencies of events from samples in the previous step to yield probabilities.
@@ 3.3 Self-annotation and fine-tuning
We compute the calibrated probability $p=c(f)$ associated with each prediction in the held-out calibration data, and choose a mapping into discrete confidence bins. Several options are possible for this binning function $b$, including adaptive schemes as well as uniform schemes, the number of bins $B$, and so on. In our experiments, we focus on a simple fixed-width scheme with $5$ bins. Let $\hat{Y}$ denote the set of all predictions on $X^{\text{cal}}$, and, if the model was previously fine-tuned on a supervised training set $X^{\text{train}}$, we include predictions on $X^{\text{train}}$. We transform each prediction and calibrated confidence into a training example for a round of supervised fine-tuning by verbalizing the corresponding bin in the answer. For example, the fifth of five bins may correspond to “very high confidence.” The token sequences chosen to encode each bin are arbitrary, as we discuss in Appendix F; for easy interpretability, we use short confidence descriptors in this paper, namely “very low,” “low,” “medium,” “high,” and “very high.”
@@
While correct answers should be added as training data, appending the confidence scores to incorrect answers may improve the model’s ability to correctly verbalize its own confidence. However, it may also decrease the accuracy of the QA model. We introduce a hyperparameter to control the number of incorrect answers added to the training data. In § B.2, we further investigate the impact of this hyperparameter.
Starting from the sampled model, we perform supervised fine-tuning on these self-annotated targets with verbalized confidences to estimate a second model capable of verbalizing its confidence. If training an instruction-tuned model, we append an additional instruction such as “Additionally state how confident you are in your answer.” to the preexisting instruction [^7] See Appendix E for details on the specific prompts used in each experiment.. If a reasoning trace has been generated during sampling, we randomly select a reasoning trace to add to the target answer from all possible options. At inference time, we obtain predictions and verbalized confidences from this new model on held-out test data. This test data has no overlap with the post-hoc calibrated training set, and can even be drawn from an entirely different dataset, as in § 6. We remark that our model incurs little additional cost at inference time, as opposed to other confidence elicitation methods which require inference-time sampling (Farquhar2024; xiong2024can).